OpenStudy (anonymous):

Solve for a in the equation 5/2a = - 25 -10 Right

OpenStudy (anonymous):

Give me one minute to check your answer :)

OpenStudy (across):

If you mean: (5/2) * a = -25, then a = -10 If you mean: 5/(2 * a) = -25, then a = 1/-10

OpenStudy (anonymous):

So i gots it right?

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

Thank you satellite

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

\[a = - 1/10\]

OpenStudy (anonymous):

correct, a=-10

OpenStudy (anonymous):

Thank you abhis

OpenStudy (anonymous):

my pleasure:)

OpenStudy (anonymous):

Your answer should be - 1/10. O.k. Please click good answer and give a medal :)

OpenStudy (anonymous):

lol l8 to the party man

OpenStudy (anonymous):

Can you click good answer so I can get a medal :)